OpenStudy (anonymous):

can you please help me solve 3x^2=-x+10 by factoring?

OpenStudy (anonymous):

To solve... it need to = 0 first 3x^2 + x - 10 = 0 now factor (3x - 5)(x + 2) = 0 x = 3/5 or x = -2